# Reference Architecture of a Quantum-Centric Supercomputer

### 11 Mar 2026 v1

@@ II-A Use Case 1: Electronic Structure Calculations
Concretely, the system architecture features loose spatial coupling and weak temporal coupling between heterogeneous quantum and HPC systems. The quantum system can either be co-located with the classical HPC system or connected over the cloud, and both systems are managed by independent workload managers operating in batch-time mode. In the SQD workflow, the chemistry Hamiltonian is first mapped to qubits, then the QPU generates bitstring samples representing electronic configurations (Slater determinants). The classical HPC supercomputer performs (1) configuration recovery to incorrect sampling artifacts (e.g., particle number errors due to noise), (2) subsampling of recovered configurations, and (3) subspace projection and diagonalization of the Hamiltonian in the selected determinant basis (see Figure 1 in  [83]). A key architectural point, shown in that figure, is that the SQD process on the classical HPC is executed in a self-consistent loop among these three steps: the diagonalization outputs update a reference occupancy vector, which is fed back into the configuration recovery and the loop repeats until convergence or exceeding a certain number of iterations. Here, the QPU provides the sampling signal and the HPC system iterates and refines the sparse wave function approximation as shown in Figure 2(a).
@@
Key Observation 1: SQD exemplifies batch-time integration with loose coupling between quantum and classical HPC systems. These workflows do not require spatial or temporal co-location of systems, enabling flexible deployment across distributed or cloud connected infrastructures.
@@ II-B Use Case 2: Closed-Loop Electronic Structure Calculations
Shirakawa et al. [89] demonstrated closed-loop electronic structure calculations using an on-premises IBM Heron Quantum system along with the Fugaku supercomputer. This work introduces a more tightly integrated orchestration model that treats the quantum-classical workflow as a co-scheduled distributed optimization problem. Instead of simply iterating an estimator loop as in use case 1, this method iteratively refines a quantum state representation of the molecular wavefunction using quantum sampling, classical subspace diagonalization, and parameter optimization as shown in Figure 2(b). The loop continues until the ground-state energy converges. This architecture enables electronic structure calculations beyond the reach of traditional classical full configuration interaction methods (e.g., CCSD, HF) by combining quantum sampling efficiency with massively parallel classical computations.
@@
A key characteristic of this workflow is the tight temporal and spatial coupling between quantum hardware and the classical HPC system. This coupling is required for a few reasons: (1) continuous data exchanges between the classical system and the quantum system as each completes their step, the output is used in the following step on the other system, (2) because the quantum and classical computations are in a loop where both system resources are essentially dedicated to this workflow, the data must be transferred in a timely manner from a latency and bandwidth perspective, and (3) It is possible to dedicate both systems for this workflow simply because both systems are co-located and managed by a single organization that has control on the job queues and user priorities on these systems. If either or both of these systems are accessed via cloud, and shared with a large group of users from different organization, dedicating time across both systems will be a significantly challenging task. For these reasons, a temporally and spatially co-located Quantum and Classical HPC system is desired to orchestrate such closed loop computations.
@@
Figure 2: Illustration of the (a). SQD workflow and (b). Closed-loop SQD workflow between Quantum and classical resources. In (a). the results of the classical post processing are used to compute ground-state energies, where as in (b). the results of the classical step are fed to optimize the quantum circuits in subsequent steps to achieve optimal ground-state energies.
@@ II-E Use Case 5: Outer code research for fault-tolerant Quantum Error Correction
Fault-tolerant quantum computing will likely rely on hierarchical error correction, in which multiple layers of coding operate at different physical and algorithmic timescales. Rather than relying on a single monolithic code, this approach separates the problem into an inner hardware-adapted code that stabilizes physical qubits in real time and an outer algorithm-adapted code that provides additional logical protection and noise suppression 3. The concept builds on the original theory of concatenated quantum codes, where logical qubits are recursively encoded to reduce logical error rates below threshold. Modern developments in quantum error correction, including recent advances in quantum low-density parity check (LDPC) codes and scalable fault-tolerant architectures further motivate layered approaches that align each level of coding with the latency and computational capabilities of surrounding classical control and compute systems.
@@ III QCSC Reference Architecture
Figure 4: Quantum Centric Supercomputing Architecture: the architecture describes how quantum systems can be integrated with high-performance computing infrastructure (CPUs + GPUs) in a composable compute layer, with system orchestration, application middleware, and applications with domain-specific libraries built on top.
@@ III-A QCSC Hardware Infrastructure Layer
The Partner Scale-up Co-located Systems, in the middle, are CPU and GPU systems from the likes of Intel, AMD, or NVIDIA that are physically in proximity to the Quantum System and connected via a low-latency interconnect such as RDMA over Converged Ethernet (ROCE) [43] , Ultra Ethernet [37], or NVQLink [16]. We call this a near-time interconnect because of the expected proximity between this scale-up system and the Quantum system as well as the low-latency interconnect that connects these systems. Here, we anticipate workloads such as in Use Case 5, with the potential for research into new decoding algorithms for outer codes within a hierarchical fault-tolerant error correction framework.
To enable this a user needs access to a programmable GPU compute resource (e.g., shown in Figure LABEL:fig:decoder). The co-located GPU and FPGA nodes are well-suited for this task such as training and evaluating decoders against live or recorded syndrome streams for the QPUs.
The Partner Scale-out Systems (CPUs+GPUs, Cloud or On-Premises) are connected to the combined Scale-up co-located systems and the Quantum Systems via an appropriately sized high-bandwidth interconnect and represent the broadest and the most flexible classical compute resources in the QCSC stack. These systems are not constrained to a single hardware configuration: they may be deployed as pure GPU clusters optimized for massively parallel HPC and AI workloads, or as mixed GPU and CPU clusters to support traditional modeling and simulation workloads, and Quantum workloads. This flexibility allows the Quantum systems to be deployed in existing HPC data centers in the near-term, or planned AI factories being scaled out for the future. This configurability also allows the Scale-out tier to be tailored to the specific classical computation profile of the quantum application domain it serves. The Scale-out tier handles the full range of classical workloads that surround and complement QPU execution, including:
@@ III-E1 QCSC Cloud Software
At the same time, broader trends in HPC are begging to mitigate this challenge for a different reason. Traditionally, HPC systems were primarily used for large-scale modeling and simulation workloads, which are inherently batch-oriented. However, HPC infrastructure is increasingly being used for AI workloads, which include not only large-scale training but also fine-grained services such as real-time inference and web-based AI applications, there has been a growing adoption of cloud-native software stacks. Supporting these workloads has led to the growing adoption of cloud-native software stacks, containerized services, and orchestration frameworks such as Kubernetes and OpenShift within HPC environments [32].

### 12 Mar 2026 v2

@@ II-A Use Case 1: Electronic Structure Calculations
Concretely, the system architecture features loose spatial coupling and weak temporal coupling between heterogeneous quantum and HPC systems. The quantum system can either be co-located with the classical HPC system or connected over the cloud, and both systems are managed by independent workload managers operating in batch-time mode.
In the SQD workflow, the chemistry Hamiltonian is first mapped to qubits, then the QPU generates bitstring samples representing electronic configurations (Slater determinants).
The classical HPC supercomputer performs the following operations on the samples (see Figure 1 in [87]):
  • Partitions the samples into two groups based on hamming weight:
    1. 1.
      Correct hamming weight: Samples with the expected number of electrons proceed unchanged to the next step
    2. 2.
      Incorrect hamming weight: Samples with incorrect hamming weights due to noise undergo configuration recovery to restore the correct hamming weight matching the number of electrons (recovered samples)
  • Combines the correct hamming weight samples with the recovered samples
  • Performs subsampling on the combined set
  • Executes subspace projection and diagonalization of the Hamiltonian in the selected determinant basis (the most computationally intensive step)
A key architectural point, shown in that figure, is that the SQD process on the classical HPC is executed in a self-consistent loop among these three steps: the diagonalization outputs update a reference occupancy vector, which is fed back into the configuration recovery and the loop repeats until convergence or exceeding a certain number of iterations. Here, the QPU provides the sampling signal and the HPC system iterates and refines the sparse wave function approximation as shown in Figure LABEL:fig:sqd(a).
@@
Key Observation 1: SQD exemplifies batch-time integration with loose coupling between quantum and classical HPC systems. These workflows do not require spatial or temporal coupling or co-location of systems, enabling flexible deployment across distributed or cloud connected infrastructures.
@@ II-B Use Case 2: Closed-Loop Electronic Structure Calculations
Shirakawa et al. [93] demonstrated closed-loop electronic structure calculations using an on-premises IBM Heron Quantum system along with the Fugaku supercomputer. This work introduces a more tightly integrated orchestration model that treats the quantum-classical workflow as a co-scheduled distributed optimization problem. Instead of simply iterating an estimator loop as in use case 1, this method iteratively refines a quantum state representation of the molecular wavefunction using quantum sampling, classical subspace diagonalization, and parameter optimization as shown in Figure LABEL:fig:sqd(b). The loop continues until the ground-state energy converges. This architecture enables electronic structure calculations beyond the reach of traditional classical full configuration interaction methods (e.g., CCSD, HF) by combining quantum sampling efficiency with massively parallel classical computations.
@@
A key characteristic of this workflow is the tight temporal and spatial coupling between quantum hardware and the classical HPC system. This coupling is required for a few reasons: (1) continuous data exchanges between the classical system and the quantum system as each completes their step, the output is used in the following step on the other system, (2) because the quantum and classical computations are in a loop where both system resources are essentially dedicated to this workflow, the data must be transferred in a timely manner from a latency and bandwidth perspective, and (3) It is possible to dedicate both systems for this workflow simply because both systems are co-located and managed by a single organization that has control on the job queues and user priorities on these systems. If either or both of these systems are accessed via cloud, and shared with a large group of users from a different organization, dedicating time across both systems will be a significantly challenging task. For these reasons, a temporally and spatially co-located Quantum and Classical HPC system is desired to orchestrate such closed loop computations.
@@ II-E Use Case 5: Outer code research for fault-tolerant Quantum Error Correction
Fault-tolerant quantum computing will likely rely on hierarchical error correction, in which multiple layers of coding operate at different physical and algorithmic timescales. Rather than relying on a single monolithic code, this approach separates the problem into an inner hardware-adapted code that stabilizes physical qubits in real time and an outer algorithm-adapted code that provides additional logical protection and noise suppression 2. The concept builds on the original theory of concatenated quantum codes, where logical qubits are recursively encoded to reduce logical error rates below threshold. Modern developments in quantum error correction, including recent advances in quantum low-density parity check (LDPC) codes and scalable fault-tolerant architectures further motivate layered approaches that align each level of coding with the latency and computational capabilities of surrounding classical control and compute systems.
@@ III QCSC Reference Architecture
Figure 3: Quantum Centric Supercomputing Architecture: the architecture describes how quantum systems can be integrated with high-performance computing infrastructure (CPUs + GPUs) and Heterogeneous Specialized Accelerators in a composable compute layer, with system orchestration, application middleware, and applications with domain-specific libraries built on top.
@@ III-A QCSC Hardware Infrastructure Layer
The Partner Scale-up Co-located Systems, in the middle, are CPU, GPU, and Accelerator systems from the likes of Intel, AMD, or NVIDIA that are physically in proximity to the Quantum Systems and connected via a low-latency interconnect such as RDMA over Converged Ethernet (ROCE) [45] , Ultra Ethernet [39], or NVQLink [16]. We call this a near-time interconnect because of the expected proximity between this scale-up system and the Quantum system as well as the low-latency interconnect that connects these systems. Next-generation scale-up systems will feature heterogeneous specialized accelerators (simply Accelerators) purpose-built for emerging artificial intelligence (AI) workloads [19]. Here, we anticipate workloads such as in Use Case 5, with the potential for research into new decoding algorithms for outer codes within a hierarchical fault-tolerant error correction framework.
To enable this a user needs access to a programmable compute resource (e.g., shown in Figure LABEL:fig:decoder). The co-located CPU, GPU, and Accelerator nodes are well-suited for this task such as training and evaluating decoders against live or recorded syndrome streams for the QPUs.
The Partner Scale-out Systems (CPUs+GPUs+Accelerators, Cloud or On-Premises) are connected to the combined Scale-up co-located systems and the Quantum Systems via an appropriately sized high-bandwidth interconnect and represent the broadest and the most flexible classical compute resources in the QCSC stack. These systems are not constrained to a single hardware configuration: they may be deployed as pure GPU clusters optimized for massively parallel HPC and AI workloads, or as mixed GPU, Accelerator, and CPU clusters to support traditional modeling and simulation workloads, and Quantum workloads. This flexibility allows the Quantum systems to be deployed in existing HPC data centers in the near-term, or planned AI factories being scaled out for the future. This configurability also allows the Scale-out tier to be tailored to the specific classical computation profile of the quantum application domain it serves. The Scale-out tier handles the full range of classical workloads that surround and complement QPU execution, including:
@@ III-E1 QCSC Cloud Software
At the same time, broader trends in HPC are beginning to mitigate this challenge for a different reason. Traditionally, HPC systems were primarily used for large-scale modeling and simulation workloads, which are inherently batch-oriented. However, HPC infrastructure is increasingly being used for AI workloads, which include not only large-scale training but also fine-grained services such as real-time inference and web-based AI applications, there has been a growing adoption of cloud-native software stacks. Supporting these workloads has led to the growing adoption of cloud-native software stacks, containerized services, and orchestration frameworks such as Kubernetes and OpenShift within HPC environments [33].